OpenStudy (anonymous):

detemine if (-2/3.1) is a solution to -9x-2x=8

OpenStudy (anonymous):

-9(-2/3.1)-2(-2/3.1)=-7.10 So, no.

OpenStudy (campbell_st):

well simplify the equation -11x = 8 substitute your value -2/3.1 so -11(-2/3.1) = 22/3.1

OpenStudy (anonymous):

ok i really need some help with graphing this equation given the linear equation y=-1/2x+5, find the y-coordinates of the points (-6),(-2), and (8)

OpenStudy (campbell_st):

ok so sunstitute x = -6, -2 and 8 y = -(-6)/2 + 5 = 8 y = -(-2)/2 + 5 = 6 y = -(8)/2 + 5 = 1 you have the ordered pairs (-6, 8), (-2, 6) and (8, 1) plot the points in the line plane and you should be able to draw one line that passes through all 3 points

OpenStudy (anonymous):

Thank you so much